# Paired Completion: Flexible Quantification of Issue-framing at Scale with LLMs

### v1

@@ 4 Evaluation Approach
We compare the novel paired completion method with a total of four comparison approaches, representing a mix of traditional NLP and transformer-based LLM methods – three use a trained logistic regression (Hosmer Jr, Lemeshow, and Sturdivant 2013) classifier over varying training sample sizes, either employing TF-IDF vectors (Sparck Jones 1972; Salton 1983), word embedding vectors (Mikolov et al. 2013), or LLM contextual embeddings (OpenAI 2024b) to represent texts in high dimensional space (see Supplementary Information for details); whilst one uses LLMs via a prompt-based approach (described below).
@@ 4.1 LLM Prompting
We use a fixed prompt, which was initially fine-tuned for GPT-4 and GPT-3.5, and then further tuned for Mixtral-8x7b-Instruct-v0.1 and LLaMA-2-70B-Chat. In hindsight, it was a mistake to tune our prompts for GPT-4 first, as while GPT-4 was very likely going to give the best performance on the tasks at hand (compared to the other models in consideration), it was also a lot more forgiving of errors, confusing wording, and conflicting instructions within the prompt. It is also possible this somewhat biased the prompts towards the OpenAI models, but this was unavoidable given our preliminary results (not included in this paper) were gathered using only the OpenAI APIs; in any case, our results demonstrate the superiority of the open-source models on these tasks. We used a single prompt across all models in our final experiments.
@@ 4.2 Performance Analysis
4.3 Summary of experiments
@@ 5.1 Comparative Analysis of Classification Methods
The three methods for interaction with LLMs that were analysed are all effective, and uniquely suited to different scenarios. The paired completion approach proved highly effective, efficient, and robust to model-specific influences. Embedding-based methods are extremely cheap due to a combination of cheap models and fewer calls to the APIs, and proved very effective with sufficient data. However, this data threshold was far beyond the five exemplars used for the other two approaches, and performance suffered greatly when using 50 exemplars (which is still an order of magnitude more than the five exemplars provided to the other two methods). Prompt-based completion proved effective, particularly with GPT-4 (which does not support the logit outputs required for paired completion), but when possible we generally found paired completion to be more cost-effective than the prompting approach. On the other hand, the prompt-based method is very straightforward, and as models increase in capability (and, ideally, increase in cost-performance as well) the cost-performance distinction between these two techniques might diminish.
@@ 5.3 Model Bias
We observed differences in the bias displayed by models and techniques that were dataset-dependent (Figure 4). Embedding-based approaches appear most robust to bias, with no statistically significant bias found for any configuration. LLM-based approaches demonstrated bias in some scenarios, with the $k=2$ paired completion configuration potentially reducing bias compared to $k=1$. The top performing LLM paired completion methods (mxtrl-k=2; llama-k=2) show significantly less bias than the top LLM prompting approaches, including GPT-4. Further studies are needed to examine the sources of these biases, such as bias in training data, language modeling, or alignment. However, the results suggest the stronger LLM paired completion methods (e.g. llama-k=2) achieve a balance of high accuracy and low bias.
@@ 6 Limitations & Further Work
- Whilst using synthetic data for evaluation has some benefits (described earlier) there are also some significant limitations. Principally, we cannot be sure that our evaluation results would carry over to a human-labelled ground-truth data set. In the ideal case, a large expert annotation activity would be undertaken to generate a conceptual framing dataset covering a range of issues, and dimensions, and conceptual framings. Such a dataset would be of huge benefit to the field and would no doubt spur further refinement of framing analysis methods. As an intermediate step, a representative sample of our synthetic dataset could be validated by expert annotation to provide some comfort to our main findings.
Opposing framings, extension beyond ‘support,oppose’
– Our study leverages a traditional two-class classification paradigm by seeking each method’s determination as to whether a given target text is drawn from synthetic framing set ‘a’ or ‘b’. Whilst parsimonious and a reasonable starting position, a natural question arises as to how paired completion could be extended to any number of framings on a given dimension of an issue. For example, climate change (issue) causation (dimension) could be explored with three framings around: ‘anthropogenic activity’, ‘volcanic activity’, or ‘solar cycles’. The time complexity of the comparison algorithm after calling the model is possibly superlinear in the number of classes (though the exact behaviour would be implementation-dependent), but the number of model calls is linear in the number of classes. We thus expect the latter to be the overwhelming factor at least until reaching thousands of classes, given an efficient implementation of the comparison algorithm.

### v2

@@ 3 Evaluation Approach
We compare the novel paired completion method with a total of four comparison approaches, representing a mix of traditional NLP and transformer-based LLM methods – three use a trained logistic regression (Hosmer Jr et al., 2013) classifier over varying training sample sizes, either employing TF-IDF vectors Sparck Jones (1972); Salton (1983), word embedding vectors Mikolov et al. (2013), or LLM contextual embeddings OpenAI (2024b) to represent texts in high dimensional space (see the appendices for details); whilst one uses LLMs via a prompt-based approach (described below).
@@ 3.1 LLM Prompting
We used a fixed prompt across all models, initially optimized for GPT-4 and GPT-3.5, then adapted for Mixtral and LLaMA-2. Starting with GPT-4 proved suboptimal, as its tolerance for prompt imperfections may have influenced our design choices. While this approach potentially favored OpenAI models, our results ultimately showed open-source models performing better on the target tasks. We used a single prompt across all models in our final experiments.
@@ 3.2 Performance Analysis
3.3 Dataset Formation
Existing large framing datasets like the MFC Card et al. (2015) and GVFC Liu et al. (2019a) focus on labeling dimensions of topics rather than conceptual framings, making them unsuitable for our task. We therefore developed a synthetic dataset generation pipeline to create balanced, controlled examples of contrasting framings across multiple topics to provide a rich evaluation setting. In a second study, we use a small, human-labelled dataset on Immigration Tweets Mendelsohn et al. (2021b) to study paired-completion in more realistic settings. Our synthetic approach uses a two-step hierarchical process: first generating seed perspectives on a given topic, then producing sentences that align with each perspective. The pipeline ensures balanced representation across linguistic features to avoid confounding effects.
While synthetic data has inherent limitations for evaluation, it provides key advantages: controlled variation of framing elements, balanced representation across perspectives, and mitigation of potential training data overlap with the LLMs being evaluated. We generated datasets for four topics of varying complexity: dog ownership (as a straightforward baseline), climate change, domestic violence, and misogyny (representing more nuanced, controversial issues). Detailed generation protocols and validation procedures are provided in the appendices.
3.4 Summary of Synthetic Experiments
@@
3.5 The Immigration Tweet Dataset
To evaluate the performance of paired-completion in a more realistic setting, we applied paired completion to the labelled dataset from Mendelsohn et. al Mendelsohn et al. (2021b) [^2] We thank the authors for providing this dataset., focusing on the “dev” dataset of $450$ high-quality labelled tweets across various categories. All tweets were consensus labelled by two trained annotators with one or more of 11 specific frames (e.g. ‘Humanitarian: Immigrants experience economic, social, and political suffering and hardships’), within three main frames (‘threat’, ‘hero’, ‘victim’), with the source containing a description of each frame (see Appendix for details). The descriptions were passed to Claude 3.5 Sonnet to create exemplar tweets as conditioners (nb: not the tweets from the dataset). To mimic the contrasting textual-alignment task with the tweet dataset, we re-sample the dataset into 55 subsets, where each subset contained a unique combination of $A|B$ labels ($c(c-1)/2$, for $c=11$ categories). The largest so formed subset contained 134 tweets (A:Humanitarian—B:Public Order), whilst the smallest contained just 9 tweets (A:Global Economy—B:War).
@@ 4.3 Model Bias
We observed differences in the bias displayed by models and techniques that were dataset-dependent (see Appendix for details). Embedding-based approaches appear most robust to bias, with no statistically significant bias found for any embedding configuration. LLM-based approaches demonstrated bias in some scenarios, with the $k=2$ paired completion configuration potentially reducing bias compared to $k=1$. The top performing LLM paired completion methods (mxtrl-k=2; llama-k=2) show significantly less bias than the top LLM prompting approaches, including GPT-4. Further studies are needed to examine the sources of these biases, such as bias in training data, language modeling, or alignment. However, the results suggest the stronger LLM paired completion methods (e.g. llama-k=2) achieve a balance of high accuracy and low bias.
4.4 Immigration Tweets
We next apply the most performant paired-completion configuration from the synthetic study to the Immigration Tweet dataset from Mendelsohn et al. (2021b).
Figure 4: Immigration Tweet evaluation Bars and error-bars indicate mean and 95% CIs for F1 scores of all experiments within a main category pairing, whilst black markers indicate individual experiment means (x-axis jitter applied).
In fig. 4 we present mean F1 scores across the 55 unique A-B experiments, grouped by main frame pairs. Averages for these groupings (reading L-R in the figure) were 0.51, 0.61, 0.57, 0.43, 0.65, 0.47 respectively. The highest F1 scores were obtained for the A:Public Order(threat)–B:National Cohesion(threat) (F1$=0.872$), A:Humanitarian(victim)–B:Cultural Diversity(hero) (F1$=0.864$), and A:Humanitarian(victim)–B:Integration(hero) (F1$=0.835$) pairs. In general, there is some indication that performance is stronger where the two frames are more distinct, i.e. taken from different main framing categories. For instance, the least peformant experiment was for A:Global Economy(victim)–B:Humanitarian(victim) (F1$=0.091$), with specific frame descriptions being barely distinct, A:‘Immigrants are victims of global poverty, underdevelopment and inequality’ vs B:‘Immigrants experience economic, social, and political suffering and hardships’. Whilst direct comparison with the reference is not possible since they train a multi-stage multi-class labelling pipeline with thousands of labelled tweets, it is of note that the reported mean F1 score (across 11 specific categories) for the Issue-specific ‘dev’ dataset is 0.550 (Table 3 in the reference), equivalent to our paired-completion findings obtained without training.
@@ 5 Limitations & Further Work
- Whilst using synthetic data for evaluation has some benefits (described earlier) there are also some significant limitations. Principally, although our results with the small Immigration Tweets dataset are promising, ideally, a large expert annotation activity should be undertaken to generate a conceptual framing dataset covering a range of issues, and dimensions. Such a dataset would be of huge benefit to the field and would no doubt spur further refinement of framing analysis methods. As an intermediate step, a representative sample of our synthetic dataset could be validated by expert annotation to provide some comfort to our main findings.
Opposing framings, extension beyond binary classification
The most performant evaluation results from the Immigration Tweets dataset back up our findings in the synthetic dataset, namely that paired completion is most effective in opposing framing contexts. However, it is possible that three or four distinct framings may also be successfully identified in corpora. The method’s computational complexity scales linearly with the number of classes in terms of model calls, making it feasible for practical multi-class applications. Tentative experiments along these lines using the main frames in the immigration dataset suggest this is the case, but it remains for future work to explore the limits of paired completion in a many-frames setting.